OpenStudy (bohotness):

@jdoe0001 @livinitup

OpenStudy (jdoe0001):

2 ducks and 1 swan

OpenStudy (anonymous):

?

OpenStudy (bohotness):

:3

OpenStudy (bohotness):

. If the diameter of the circle above is 12 mm, what is the area of the circle? A. 144 mm2 B. 6 mm2 C. 12 mm2 D. 36 mm2

OpenStudy (jdoe0001):

well... that's the answer, unless you have an specific :)

OpenStudy (bohotness):

 a ?

OpenStudy (bohotness):

nvm

OpenStudy (anonymous):



OpenStudy (anonymous):

FIRST STEP:FIIND THE RADIUS

OpenStudy (jdoe0001):

\(\bf area = \pi r^2\qquad diameter =12\qquad r=\frac{diameter}{2} \\ \quad \\ area = ?\)

OpenStudy (anonymous):

second: multiply by 3.14

OpenStudy (bohotness):

6

OpenStudy (anonymous):

third: put it to the 2nd power

OpenStudy (bohotness):

its is 6?

OpenStudy (jdoe0001):

\(\bf area = \pi r^2\qquad diameter =12\qquad r=\frac{diameter}{2}=\frac{\cancel{ 12}}{\cancel{ 2}}=6 \\ \quad \\ area = \pi\cdot 6^2\)

OpenStudy (bohotness):

so its 6?

OpenStudy (anonymous):

yes 6 is the radius

OpenStudy (anonymous):

now you have to multiply it by 3.14 or pi

OpenStudy (anonymous):

its not the real answer

OpenStudy (bohotness):

wyhatttttt

OpenStudy (anonymous):

multiply 6 by 3.14

OpenStudy (bohotness):

O.OOOOOOOOO wahat i throughti t was g

OpenStudy (anonymous):

you can use a calculator

OpenStudy (jdoe0001):

ahemm, "someone", no name sorta left out that I belive the answer is meant to be "in terms of \(\Large \pi\)"

OpenStudy (bohotness):

you answer i got is not my ansewr CHOICE

OpenStudy (jdoe0001):

\(\bf area = \pi r^2\qquad diameter =12\qquad r=\frac{diameter}{2}=\frac{\cancel{ 12}}{\cancel{ 2}}=6 \\ \quad \\ area = \pi\cdot 6^2\qquad {\color{brown}{ 6^2=36}}\qquad area = {\color{brown}{ 36}}\pi\)

OpenStudy (bohotness):

If the radius of the circle above is 10 mm, what is the circumference of the circle? A. 20 mm B. 40 mm C. 100 mm D. 10 mm okay

OpenStudy (jdoe0001):

same as the one you did before this one use the same equation to get the circumference \(\bf c = 2\pi r\)

OpenStudy (bohotness):

so 10 times 2 ? times 3.144

OpenStudy (jdoe0001):

well. you see your choices seem to imply that is in \(\Large \pi\) terms no in decimal terms so you have to keep the \(\Large \pi\) in the answer as well

OpenStudy (bohotness):

okay

OpenStudy (bohotness):

thats 62.8

OpenStudy (bohotness):

@jdoe0001

OpenStudy (jdoe0001):

well.. if the 10, \(\bf 2\pi r = 2\pi \cdot 10\) and that would give you 62.8 BUT in decimal terms which is NOT any of the choices thus, the likelyhood, check your paper, the answer is in \(\pi\) terms

OpenStudy (bohotness):

what ijust did that

OpenStudy (bohotness):

@jdoe0001